Circular No. 5723

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993d) Jean Mueller, Palomar Mountain Observatory, reports her discovery

of a comet on a 75-min exposure obtained on Mar. 19 by J. D. Mendenhall and herself with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. The comet had a very obvious tail to the northwest. A confirmatory plate was taken on Mar. 24 by Mueller and C. Brewer through clouds. Mueller has measured the following positions:

1993 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 19.35903  13^{h}01^{m}45.^{s}3  +64^{o}20′47′′  17
19.38507  13 01 44.55  +64 20 49.4
19.41112  13 01 43.92  +64 20 50.5
24.35938  12 58 12.31  +64 22 47.3

SUPERNOVA 1993H IN ESO 445-G66 M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report that M. Wischnjewsky has discovered a supernova 1′′ east and 12′′ north of the nucleus of the spiral galaxy ESO 445-G66 (α = 13^{h}49^{m}58^{s}, δ = −30^{o}27′.9, equinox 1950.0). The supernova, of m_{pg} ∼ 16.5, was found on a 20-min unfiltered IIa-O plate taken by G. Valladares and R. Antezana with the CTIO Curtis Schmidt telescope on Mar. 18.20 UT. Confirmation was made by D. Geisler from CCD R and I images obtained on Mar. 21.275 with the CTIO 0.9-m telescope. Spectra obtained by M. M. Phillips with the CTIO 4-m telescope on Mar. 21.22 and by M. Hamuy with the CTIO 1.5-m telescope on Mar. 23.13 reveal that the supernova is of type Ia. The second spectrum is intermediate in appearance between the maximum-light spectra of the type Ia supernovae 1992A and 1986G. The minimum of the Si II (635.5 nm) line yields an expansion velocity of 10 100 km/s with respect to the galaxy’s rest frame.

U GEMINORUM Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany:

Mar. 19.88 UT, 9.7; 20.86, 9.3.

Corrigendum. On IAUC 5722, for Kibbutz Sheffagim read Kibbutz Sheffayim.

1993 March 24 Brian G. Marsden